In 3d^4 , what s the sum of angular momemtum , spherical nodes and angular nodes in terms of h and pie

In 3d^4  , what s the sum of angular momemtum , spherical nodes and  angular nodes in terms of h and pie

Grade:

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please ask complete question

Regards
meghana p
3 years ago
angular momemtum = root under [l][l+1] *h\2pie=root under 6 *h\2pie
spherical nodes =3-2-1=0
angular nodes =2 for d orbital 
their sum = root6 +4pie\2pie